Find the mean, variance and standard deviation of the binomial distribution with the given values of n and p.
N = 70, P = 0.4

Answer by jim_thompson5910(35256)

Mean = np = 70*0.4 = 28

Variance = np(1-p) = 70*0.4*(1-0.4) = 16.8

Standard deviation = sqrt(Variance) = sqrt(16.8) = 4.09878